What is the slope of the line: 2y + 8x = 26
Answer by stanbon(75887)

What is the slope of the line: 2y + 8x = 26
-----
2y + 8x = 26
y + 4x = 13
y = -4x+13
-------
Slope is always the amount that y changes when x changes 1.
---
slope = -4
================
Cheers,
Stan H.